OSG Area Coordinators Meeting

Meeting Coordinates

Date Thursday, September 9th, 2010
Time Noon Central

1. Dial Toll-Free Number: 866-740-1260 (U.S. & Canada)

2. Participants enter your access code: 8405618, followed by #

Attendees

Robert Engel, Igor Sifligoi, Mine Altunay, Kent Blackburn, John McGee?, Brian Bockelman, Alain Roy, Chander Sehgal

Agenda

  • 1.4 VOs Group Abhishek Rana

  • 1.5 Engagement - John McGee? (confirmed)

  • 1.7 Security Mine Altunay (confirmed)

  • 1.8 Content Management - Robert Engel (confirmed)

1.5 Engagement

  • Steffen Bass: Met with Steffen on Aug-20. Attempted to get usage projections, but he is not willing to make an educated guess until we get further into the school year. He does expect usage ramp up at some level this fall. Steve Cox is working to better understand the security model of iRODS in order to prepare a suggested deployment scenario. Talked with David Skinner at NERSC, who suggested that the OSG staff at NERSC would be a good starting point for a test deployment. Leesa Brieger of RENCI and the iRODS team will be attending the upcoming OSG Storage Forum. Leesa, Steve, and I met to discuss what we are trying to accomplish with NERSC for Steffen. Next Steps: gather iRODS seurity model and HPSS interface expertise for NERSC, request a test deployment.

  • Have been working with an Amber user that was referred to us by SBGrid on HTPC

  • RENCI Bellhaven resource is currently down, looking into it, likely hardware failue

  • Still working on RENCI Blue Ridge resource (will be a strong HTPC resource)

  • Discussed possible new OSG CE within NCSU's VCL (cloud like environment). After RENCI CEs are operational, Steve Cox will explore deploying a test CE within the VCL that can come and go depending on available resources (provisioned or deprovisioned under VCL management).

  • General Engage VO usage has been typical, recent activity chart attached

1.7 Security

We last reported on May 20th. Ongoing items back then was:
  • IGTF new layout testing. Completed.
  • We were waiting for official pakiti release for installing the service. Completed.
  • OSG specific certificate request web page. this has been requested at All hands meeting. We tested the page's usability with CMS users center, Burt, Hover and several other scientists (approximately 20+ people). Based on the good feedback, the page is completed and tested in ITB. It is ready to be in production on 9/14
  • Concerns we reported on May 20th:
    • GUMS development. See below.
    • SBgrid visit for identity management solution. Completed the visit, reduced the registration workflow to couple steps for the end user (was 8-step).
Since the OSG staff retreat, we focused on the below items.
    • Documentation: see below

  • Identity management. We made a trip to BNL in order to gather Atlas security requirements concerning id management. John Hover and Michael Ernst went through last year's requirements documents. They are pleased with the new cert request web page and requested that it be put into production as soon as possible. We will also add a renewal feature to this page so that OSG users would not have to use DOEgrids web site. they said these two items will satisfy their urgent requirements.
    • Their longer term ID needs is having a federated Identity provider . Currently CERN provides such a service to all CERN-affiliates. CMS is also working to provide a similar service just for us-cms users. Us-atlas requested help in understanding how to provide such a service for their users
    • A small business, Galois, has won a SBIR funding from DOE to deliver identity management solutions to OSG. last month, they contacted Ruth and myself and sent their proposal. I had a meeting with them this week. they are not clear on what they plan to do except that they intend to deliver products to OSG to help in id mgmt area. They asked us to send requirements and needs. OSG's involvement will be limited to feedback and advise over their solutions.
    • We contacted CMS' identity architecture about CMS plans on having a single sign on service. Lassi will be at fermilab next week. At this stage we are gathering information and providing feedback when asked. We have no further involvement with CMS' work in this area.
  • CILogon. We started the work with Clemson University IdP?. We reviewed their practices for identity vetting. We worked with CILogon CA to securely allow these credentials in OSG. the 80% of the work in this area is complete. CILogon agreed to create a special signing policy for OSG where only trusted IdPs? will be included.
  • GUMS. Developer visited us at FNAL in June. I worked with him for a week to debug and get the code working. I met with the GUMS developer while at BNL in August again. He had not made any progress with open bug reports since June; his effort for BlueGene? project had doubled. John will hire a new developer for GUMS. Alain and I set Sept 30 as the deadline to fixing all bugs in GUMS and making a clean release. We are in constant check with John and the developer. John wants to have someone new by November.
  • Pakiti service has been functioning for the past three weeks. We monitor only CMS tier 3 sites. our goal is to get the site admins themselves to register and monitor themselves. We will monitor only as a backup person. We found critical vulnerabilities with some sites and emailed them to patch those.
  • Documentation. We invited Rob to our weekly meetings. he and Jim held a tutorial meeting, where they taught us how to review and how to make changes. Since then we made steady progress. We are on target to meet our October deadline to finish this project.
  • We are starting a new cycle of ST&E checks. We start with GOC since they already been audited by Indiana security officer. We will get the report results and then decide if the results meet OSG requirements.
  • We received a request to assess the risk introduced by service certificates (the UCSD data transfer problem) We developed a policy on using service certificates in OSG.

1.8 Content Management

We last reported on the project status during the OSG Staff Retreat in July 2010.

Document Status

Project Status

All documentation areas but one continue to increase the quality of documents using the Document Process?. The documentation area coordinators and the documentation architect meet on a weekly basis to report on the progress and discuss any issues regarding documentation. The documentation architect joins individual project telecons (Security, Storage, VO Forum, ITB) on an as-needed basis to discuss documentation work with owners, reviewers and testers for documentation.

The Operations team recently replaced the hardware hosting the Twiki. Our benchmarks show that this increased the performance by a factor of 2 to 3 already. Scott Teige further suggested to test using a ramdisk hosting the Twiki data outside of a virtual machine. This is work in progress for September 2010 and is expected to increase the performance not unlike a SSD without additional costs.

The twiki-docteam was setup where I started to work on navigation. I got access to the main CSS style sheet and all LeftBars? used for navigation. I reviewed the navigation concepts developed for each documentation area during the workshop in February and started to work on navigation on twiki-docteam.

The Operations area was recently created. Elizabeth has identified 39 relevant documents and integrated them into the document process. She is currently the documentation area coordinator for Operations. We are looking forward for Elizabeth to join the Content Management Project by October 2010.

The Security area made significant progress. Jim prepared a quick start tutorial which was well received by the members of the Security team. Documents are currently reviewed and some have been released.

The Compute Element area is now coordinated by Iwona and made significant process ever since. She found document owners for all pending documents and is pushing the review and test cycle on each of them.

The User area was coordinated by Derek. During his role as Summer Intern he helped progress the area to become the first area to complete the document process for all its documents. We are looking forward for Brian to replace Derek starting September 2010.

The Virtual Organization area made no progress at all since the OSG Staff Retreat and is a major concern of mine. The documentation area coordinator is practically not existent for this area. The contribution has been minimal and unsatisfactory from the very beginning. Besides repeated promises from Abhishek no progress has been made. I have difficulties contacting him by phone and e-mail and have failed to setup a meeting that would resolve the issue.

Thoughts

A Collaborative Document Rating mechanism may be as useful to increase the quality of documentation as the document process already in place. The best person to make this happen would be Tim Cartwright with 1/10 FTE. Without a dedicated effort we will not be able to support document ratings.

Who could replace Abhishek as documentation area coordinator for the VO area? I suggest to include Marcia in finding a solution to this problem.

-- JemiseLockhart - 27 Aug 2010

Topic attachments
I Attachment Action Size Date Who Comment
pngpng user_hours_last_week.png manage 49.3 K 09 Sep 2010 - 17:18 JohnMcGee Recent Engage VO usage
Topic revision: r12 - 09 Sep 2010 - 18:22:52 - ChanderSehgal
Hello, TWikiGuest
Register

 
TWIKI.NET

TWiki | Report Bugs | Privacy Policy

This site is powered by the TWiki collaboration platformCopyright by the contributing authors. All material on this collaboration platform is the property of the contributing authors..